OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (1, –2) and has a slope of -3 y = x +

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

did you copy and paste this question without looking at it maybe?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

Well, the equation is y = mx + b, where m is the slope, and b is the y-intercept.

OpenStudy (misty1212):

ooh now we see the slope

OpenStudy (anonymous):

First we plug it into point-slope form: \(y - y_1 = m(x - x_1)\) Where y1 is the y-value of the point, 'm' is the slope, and x1 is the x-value of the point.

OpenStudy (anonymous):

So in this case: \(y_1 = -2\) \(x_1 = 1\) \(m = -3\) Can you plug this into point-slope form? \(y - y_1 = m(x - x_1)\)

OpenStudy (misty1212):

no

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

We first put it in point-slope form, then we simplify to make it in slope-intercept form.

OpenStudy (anonymous):

We distribute the slope into the parenthesis, and add/subtract 'y1' to both sides..

OpenStudy (anonymous):

y+2=-3(x-1) ?

OpenStudy (anonymous):

Yep, you got it.

OpenStudy (misty1212):

yes that looks good

OpenStudy (anonymous):

So now we distribute -3 into the parenthesis: \(y + 2 = -3(x - 1)\) \(y + 2 = -3x + 3\) Now subtract 2 to both sides, what do you get?

OpenStudy (anonymous):

-3x+1

OpenStudy (anonymous):

Yep, so our final answer is: \(y = -3x + 1\)